OpenStudy (anonymous):

a machine costing $1400 depreciates in value 10% each year. What will it be worth 4 years from now?

OpenStudy (anonymous):

retriceumption is straightline depreciation 1400 * .10 = 140 depreciate annually * 4 = 560 1400 - 560 = 840

OpenStudy (anonymous):

sorry * assumption is based on straightline depreciation

OpenStudy (anonymous):

1st year $1400*(0.10)= 140 2nd year 1400-140 = 1260 * 0.10 = 126 3rd 1260-126 = 1134*0.10 = 113.4 4th 1134-113.4 = 1020.6*0.10 = 102.06 1020.6-102.06 = $918.54 after 4 years

OpenStudy (anonymous):

actually, depreciation calculations do not get adjusted like this annually. this is set when equipment is initially placed in service. So annually, depreciation for this eequipment is $140. there are differeent methods, straighline vs double declining - but NONE takes calculations like you posted. I have a MBA in accounting and finance. :)